# Moments of characteristic polynomials and their derivatives for $SO(2N)$ and $USp(2N)$ and their application to one-level density in families of elliptic curve $L$-functions

### v1

@@ 2.1.6 Pulling together the case of one pole away from zero
×$\times\left[\mathcal{M}^{SO(2N)}+\frac{\mathcal{J}^{SO(2N)}}{N}\left(z(-\alpha)-z(-\gamma)+\frac{K-2}{2}\right)\right]$
@@
×$\times\ \mathcal{M}^{SO(2N)}\left(1+\frac{(K-1)(K-2)}{2N}\left(z(-\alpha)-z(-\gamma)+\frac{K-2}{2}\right)\right),$ (128)
@@
The corresponding contribution from a pole at $-\alpha$ is simply
@@
×$\times\mathcal{M}^{SO(2N)}\left(1+\frac{(K-1)(K-2)}{2N}\left(z(\alpha)-z(-\gamma)+\frac{K-2}{2}\right)\right).$ (129)
@@
×$\times\left[\mathrm{e}^{N\alpha}\frac{z(\alpha)^{K-1}}{z(\alpha+\gamma)}\left(1+\frac{(K-1)(K-2)}{2N}\left(z(-\alpha)-z(-\gamma)+\frac{K-2}{2}\right)\right)\right.$
@@
×$\times\oint\dots\oint\frac{\displaystyle\prod_{1\leq j<k\leq K-1}z\left(w_{j}+w_{k}\right)\exp\left(N\sum_{l=1}^{K-1}w_{l}\right)\Delta(w_{1}^{2},w_{2}^{2},\dots,w_{K-1}^{2})^{2}\displaystyle\prod_{i=1}^{K-1}w_{i}\mathrm{d}w_{i}}{\displaystyle\prod_{n=1}^{K-1}w_{n}^{2K-2}},$ (131)
@@
×$\qquad\times N^{(K-1)(K-2)/2}\left(\mathcal{M}^{SO(2N)}+\frac{K-2}{2N}\mathcal{J}^{SO(2N)}+\mathcal{O}\left(\frac{1}{N^{2}}\right)\right).$
@@ 2.2 Completing the mixed moment for $SO(2N)$
Returning to (2), we have found that in the sum over the $\epsilon$’s the leading and next-to-leading order terms when $N$ is large result from one variable being integrated around a pole at $\alpha$ or $-\alpha$ with all the other variables integrated around the pole at zero. The leading order term is of order $N^{(K-1)(K-2)/2}$ and when all poles are evaluated around zero the result is at most order $N^{\tfrac{1}{2}K^{2}-\tfrac{5}{2}K-1}$ so this is at least two orders lower than the leading order as long as $K\geq 0$.
@@
Plugging these into (2) and replacing $K-1$ with $r$, we have
@@
=\ = $2^{r^{2}/2}N^{r(r-1)/2}\left(1+\mathcal{O}\left(N^{-2}\right)\right)$
×$\times\left[\frac{z(2\gamma)z(\alpha)^{r}}{z(\alpha+\gamma)z(\gamma)^{r}}\left(1+\frac{r(r-1)}{2N}\left(z(-\alpha)-z(-\gamma)+\frac{r-1}{2}\right)\right)\right.$
@@
×$\times\frac{G(r+1)\sqrt{\Gamma(2r+1)}}{\sqrt{G(2r+1)\Gamma(r+1)}}.$ (137)
To get the logarithmic derivative, we first differentiate with respect to $\alpha$, noting that $z^{\prime}(x)=z(x)z(-x)$, and that $\mathrm{d}/\mathrm{d}x\left[z(x)^{-1}\right]=\mathrm{e}^{-x}$. Then, we set $\alpha=\gamma=\phi$: :
@@
×$\times\left[\left(rz(-\phi)-z(-2\phi)\right)\left(1+\frac{r(r-1)^{2}}{2N}\right)-z(\phi)z(-\phi)\frac{r(r-1)}{2N}\right.$
@@ 3 A mixed moment for $USp(2N)$
×$\times\oint\frac{\displaystyle\prod_{1\leq j\leq k\leq K}z(w_{j}+w_{k})\exp\left(N\displaystyle\sum_{l=1}^{K}w_{l}\right)\Delta(w_{1}^{2},\dots,w_{K}^{2})^{2}\prod_{i=1}^{K}w_{i}\mathrm{d}w_{i}}{\displaystyle\prod_{n=1}^{K}z(w_{n}+\gamma)(w_{n}-\alpha)(w_{n}+\alpha)w_{n}^{2K-2}},$ (139)
where $z(x)=(1-\mathrm{e}^{-x})^{-1}=\left(\frac{1}{x}+\frac{1}{2}+\mathcal{O}\left(x\right)\right)$, and the contours of integration enclose 0, $\alpha$ and $-\alpha$. Following the steps in Section 2 exactly, one finds
$\int_{USp(2N)}-\mathrm{e}^{-\phi}\frac{\Lambda(1)^{r}{\Lambda^{\prime}(\mathrm{e}^{-\phi})}}{\Lambda(\mathrm{e}^{-\phi})}\mathrm{dA}$
@@
×$\times\left[\left(2z(2\phi)z(-2\phi)+rz(-\phi)-z(-2\phi)\right)\left(1+\frac{r(r-1)(r+1)}{2N}\right)\right.$
@@ 4 Analytic continuation in $r$
Here, $C$ is a contour surrounding the segment of the real axis between $-\pi$ and $\pi$, and we have exploited the fact that all the eigenvalues of an orthogonal matrix lie on the unit circle in the complex plane. To continue, we first switch the order of integration:
@@
[…]
@@
$\ R_{f}^{T_{\chi}}=$ $\frac{1}{2\pi}\left[\int_{\pi}^{-\pi}\int_{SO(2N)}H(\log\Lambda(1)-\chi)\frac{\mathrm{e}^{\mathrm{i}(\phi+\mathrm{i}\epsilon)}\Lambda^{\prime}(\mathrm{e}^{\mathrm{i}(\phi+\mathrm{i}\epsilon)})}{\Lambda(\mathrm{e}^{\mathrm{i}(\phi+\mathrm{i}\epsilon)})}f(\phi+\mathrm{i}\epsilon)\,\mathrm{dA}\,\mathrm{d}\phi\ \right.$
@@
×$\ \times\left(2N-\mathrm{e}^{-\mathrm{i}(\phi-\mathrm{i}\epsilon)}\frac{\Lambda^{\prime}(\mathrm{e}^{-\mathrm{i}(\phi-\mathrm{i}\epsilon)})}{\Lambda(\mathrm{e}^{-\mathrm{i}(\phi-\mathrm{i}\epsilon)})}-\mathrm{e}^{-\mathrm{i}(\phi-\mathrm{i}\epsilon)}\frac{\Lambda^{\prime}(\mathrm{e}^{-\mathrm{i}(\phi-\mathrm{i}\epsilon)})}{\Lambda(\mathrm{e}^{-\mathrm{i}(\phi-\mathrm{i}\epsilon)})}\right)\,\mathrm{d}r\mathrm{dA}\,\mathrm{d}\phi$
@@
×$\times\left(2N-2\mathrm{e}^{-\mathrm{i}(\phi-\mathrm{i}\epsilon)}\frac{\Lambda^{\prime}(\mathrm{e}^{-\mathrm{i}(\phi-\mathrm{i}\epsilon)})}{\Lambda(\mathrm{e}^{-\mathrm{i}(\phi-\mathrm{i}\epsilon)})}\right)\,\mathrm{d}r\mathrm{dA}\,\mathrm{d}\phi$
@@
×$\times\left(2N\int_{SO(2N)}\Lambda(1)^{r}\mathrm{dA}-2\int_{SO(2N)}\mathrm{e}^{-\mathrm{i}(\phi-\mathrm{i}\epsilon)}\frac{\Lambda(1)^{r}\Lambda^{\prime}(\mathrm{e}^{-\mathrm{i}(\phi-\mathrm{i}\epsilon)})}{\Lambda(\mathrm{e}^{-\mathrm{i}(\phi-\mathrm{i}\epsilon)})}\mathrm{dA}\right)\,\,\mathrm{d}r\mathrm{d}\phi.$ (155)
@@
×$\times\left(2N\left(1+\frac{r(r-1)^{2}}{2N}+\mathcal{O}(N^{-2})\right)+2\mathcal{U}(N,r,\mathrm{i}(\phi-\mathrm{i}\epsilon))\left(1+\mathcal{O}(N^{-2})\right)\right)\,\,\mathrm{d}r\mathrm{d}\phi.$ (156)
@@
×$\times\left(2N\left(1+\frac{r(r-1)^{2}}{2N}\right)+2\mathcal{U}(N,r,\mathrm{i}\phi)\right)\left(1+\mathcal{O}\left(N^{-2}\right)\right)\,\,\mathrm{d}rf(\phi)\mathrm{d}\phi.$ (157)
@@
$\frac{1}{2\pi}\int_{-\pi}^{\pi}\left[\text{Residue}\left|\frac{\mathrm{e}^{-\chi r}}{r}\mathcal{V}(N,r)\right.\right.$
×$\left.\left.\quad\quad\quad\quad\times\left(2N\left(1+\frac{r(r-1)^{2}}{2N}\right)+2\mathcal{U}(N,r,\mathrm{i}\phi)\right)\,\,\right|_{r=0}\right]f(\phi)\mathrm{d}\phi$
@@
=\ = $\frac{1}{2\pi}\int_{-\pi}^{\pi}\left[2N-2z(-2\mathrm{i}\phi)-2\mathrm{e}^{-2N\mathrm{i}\phi}z(2\mathrm{i}\phi)\right]f(\phi)\mathrm{d}\phi$
=\ = $\frac{1}{2\pi}\int_{-\pi}^{\pi}\left[2N-\frac{2}{1-\mathrm{e}^{2\mathrm{i}\phi}}-2\frac{\mathrm{e}^{-2N\mathrm{i}\phi}}{1-\mathrm{e}^{-2\mathrm{i}\phi}}\right]f(\phi)\mathrm{d}\phi$
=\ = $\frac{1}{2\pi}\int_{-\pi}^{\pi}\left[2N-\frac{2(\cos(\phi)-\mathrm{i}\sin(\phi))}{-2\mathrm{i}\sin(\phi)}+\right.$
@@
=\ = $\frac{1}{2\pi}\int_{-\pi}^{\pi}\left(2N-1+\sin\left((2N-1)\phi\right)/\sin(\phi)\right)f(\phi)\mathrm{d}\phi.$ (159)
@@
The $G$-function $G(2r+1)$ has zeros at $r\in\{-n/2\|n\in\mathbb{N}\}$, each one of multiplicity $n$. If we expand $G(2r+1)^{-1}$ as a Laurent series about $-n/2$, then, we get terms proportional to $(r+n/2)^{-n},(r+n/2)^{-n+1},(r+n/2)^{-n+2},\dots$ etc. So, when we expand $G(2r+1)^{-1/2}$ around $-n/2$, we get a Puiseaux series with terms proportional to $(r+n/2)^{-n/2},(r+n/2)^{-n/2+1},(r+n/2)^{-n/2+2}\dots$ etc. Note that we do not get powers of $(r+n/2)^{-n/2+1/2},\ (r+n/2)^{-n/2+3/2}$ etc., since this series must square to the Laurent series for $G(2r+1)^{-1}$.
We also know that $\Gamma(2r+1)$ has poles at $\{-n/2\|n\in\mathbb{N}\}$. In the Laurent expansion of $\Gamma(2r+1)$ around such a point $-n/2$, there are terms proportional to $(r+n/2)^{-1},(r+n/2)^{0},(r+n/2)^{1}$ etc. This means that in the Puiseaux expansion of $\Gamma(2r+1)^{1/2}$ around the pole at $-n/2$, we have terms proportional to $(r+n/2)^{-1/2},(r+n/2)^{1/2},(r+n/2)^{3/2}$ etc. Again, we do not get integer powers in this series, since it must square to the Laurent series for $\Gamma(2r+1)$.
Consider $\Gamma(2r+1)^{1/2}G(2r+1)^{-1/2}$ at a negative integer $\{-m\|m\in\mathbb{N}\}$. The series expansion of this product at $-m$ will have terms proportional to $(r+m)^{-m-1/2},(r+m)^{-m+1/2},\dots$ etc, from multiplying the Puiseaux series together. This series will not have a $(r+m)^{-1}$ term, so will not contribute a residue to (158).
@@
Figure 1: One level density for the excised $SO(24)$ ensemble with $\chi=\log(0.0001)$, generated numerically using $10^{7}$ matrices, with the first four terms (up to $k=3$) from (158) plotted in green. Both plots are normalised to have unit area under the curve.

### v2

@@ 2.1.6 Pulling together the case of one pole away from zero
$\times\left[\mathcal{M}^{SO(2N)}+\frac{\mathcal{J}^{SO(2N)}}{N}\left(z(-\alpha)-z(-\gamma)+\frac{K-2}{2}\right)\right]$
@@
$\times\ \mathcal{M}^{SO(2N)}\left(1+\frac{(K-1)(K-2)}{2N}\left(z(-\alpha)-z(-\gamma)+\frac{K-2}{2}\right)\right),$ (128)
@@
The corresponding contribution from a pole at $\alpha$ is simply
@@
$\times\mathcal{M}^{SO(2N)}\left(1+\frac{(K-1)(K-2)}{2N}\left(z(\alpha)-z(-\gamma)+\frac{K-2}{2}\right)\right).$ (129)
@@
$\times\left[\mathrm{e}^{N\alpha}\frac{z(\alpha)^{K-1}}{z(\alpha+\gamma)}\left(1+\frac{(K-1)(K-2)}{2N}\left(z(-\alpha)-z(-\gamma)+\frac{K-2}{2}\right)\right)\right$.
@@
$\times\oint\dots\oint\frac{\displaystyle\prod_{1\leq j<k\leq K-1}z\left(w_{j}+w_{k}\right)\exp\left(N\sum_{l=1}^{K-1}w_{l}\right)\Delta(w_{1}^{2},w_{2}^{2},\dots,w_{K-1}^{2})^{2}\displaystyle\prod_{i=1}^{K-1}w_{i}\mathrm{d}w_{i}}{\displaystyle\prod_{n=1}^{K-1}w_{n}^{2K-2}},$ (131)
@@
$\qquad\times N^{(K-1)(K-2)/2}\left(\mathcal{M}^{SO(2N)}+\frac{K-2}{2N}\mathcal{J}^{SO(2N)}+\mathcal{O}\left(\frac{1}{N^{2}}\right)\right).$
@@ 2.2 Completing the mixed moment for $SO(2N)$
Returning to (2), we have found that in the sum over the $\epsilon$’s the leading and next-to-leading order terms when $N$ is large result from one variable being integrated around a pole at $\alpha$ or $\alpha$ with all the other variables integrated around the pole at zero. The leading order term is of order $N^{(K-1)(K-2)/2}$ and when all poles are evaluated around zero the result is at most order $N^{\tfrac{1}{2}K^{2}-\tfrac{5}{2}K-1}$ so this is at least two orders lower than the leading order as long as $K\geq 0$.
@@
Plugging these into (2) and replacing K1K-1 with $r$, we have
@@
$=\$$2^{r^{2}/2}N^{r(r-1)/2}\left(1+\mathcal{O}\left(N^{-2}\right)\right)$
$\times\left[\frac{z(2\gamma)z(\alpha)^{r}}{z(\alpha+\gamma)z(\gamma)^{r}}\left(1+\frac{r(r-1)}{2N}\left(z(-\alpha)-z(-\gamma)+\frac{r-1}{2}\right)\right)\right$.
@@
$\times\frac{G(r+1)\sqrt{\Gamma(2r+1)}}{\sqrt{G(2r+1)\Gamma(r+1)}}.$ (137)
To get the logarithmic derivative, we first differentiate with respect to $\alpha$, noting that $z^{\prime}(x)=z(x)z(-x)$, and that $\mathrm{d}/\mathrm{d}x\left[z(x)^{-1}\right]=\mathrm{e}^{-x}$. Then, we set $\alpha=\gamma=\phi$, where ${\rm Re}(\phi)>0$:
@@
$\times\left[\left(rz(-\phi)-z(-2\phi)\right)\left(1+\frac{r(r-1)^{2}}{2N}\right)-z(\phi)z(-\phi)\frac{r(r-1)}{2N}\right$.
@@ 3 A mixed moment for $USp(2N)$
$\times\oint\frac{\displaystyle\prod_{1\leq j\leq k\leq K}z(w_{j}+w_{k})\exp\left(N\displaystyle\sum_{l=1}^{K}w_{l}\right)\Delta(w_{1}^{2},\dots,w_{K}^{2})^{2}\prod_{i=1}^{K}w_{i}\mathrm{d}w_{i}}{\displaystyle\prod_{n=1}^{K}z(w_{n}+\gamma)(w_{n}-\alpha)(w_{n}+\alpha)w_{n}^{2K-2}},$ (139)
where $z(x)=(1-\mathrm{e}^{-x})^{-1}=\left(\frac{1}{x}+\frac{1}{2}+\mathcal{O}\left(x\right)\right)$, ${\rm Re}(\gamma)>0$, and the contours of integration enclose 0, $\alpha$ and $\alpha$. Following the steps in Section 2 exactly, one finds
$\int_{USp(2N)}-\mathrm{e}^{-\phi}\frac{\Lambda(1)^{r}{\Lambda^{\prime}(\mathrm{e}^{-\phi})}}{\Lambda(\mathrm{e}^{-\phi})}\mathrm{dA}$ (140)
@@
$\times\left[\left(2z(2\phi)z(-2\phi)+rz(-\phi)-z(-2\phi)\right)\left(1+\frac{r(r-1)(r+1)}{2N}\right)\right$.
@@ 4 Analytic continuation in $r$
Here, $C$ is a contour surrounding the segment of the real axis between $\pi$ and $\pi$, and we have exploited the fact that all the eigenvalues of an orthogonal matrix lie on the unit circle in the complex plane. To continue, we first switch the order of integration:
@@
Then, we note that the measure $\mathrm{d}A_{T_{\chi}}$ on $T_{\chi}$ is equal to dAA (Haar measure on $SO(2N)$) multiplied by a Heaviside function:
@@
$\ R_{f}^{T_{\chi}}=\$$\frac{1}{2\pi}\left[\int_{\pi}^{-\pi}\int_{SO(2N)}H(\log\Lambda(1)-\chi)\frac{\mathrm{e}^{\mathrm{i}(\phi+\mathrm{i}\epsilon)}\Lambda^{\prime}(\mathrm{e}^{\mathrm{i}(\phi+\mathrm{i}\epsilon)})}{\Lambda(\mathrm{e}^{\mathrm{i}(\phi+\mathrm{i}\epsilon)})}f(\phi+\mathrm{i}\epsilon)\,\mathrm{dA}\,\mathrm{d}\phi\ \right$.
@@
$\ \times\left(2N-\mathrm{e}^{-\mathrm{i}(\phi-\mathrm{i}\epsilon)}\frac{\Lambda^{\prime}(\mathrm{e}^{-\mathrm{i}(\phi-\mathrm{i}\epsilon)})}{\Lambda(\mathrm{e}^{-\mathrm{i}(\phi-\mathrm{i}\epsilon)})}-\mathrm{e}^{-\mathrm{i}(\phi-\mathrm{i}\epsilon)}\frac{\Lambda^{\prime}(\mathrm{e}^{-\mathrm{i}(\phi-\mathrm{i}\epsilon)})}{\Lambda(\mathrm{e}^{-\mathrm{i}(\phi-\mathrm{i}\epsilon)})}\right)\,\mathrm{d}r\mathrm{dA}\,\mathrm{d}\phi$
@@
$\times\left(2N-2\mathrm{e}^{-\mathrm{i}(\phi-\mathrm{i}\epsilon)}\frac{\Lambda^{\prime}(\mathrm{e}^{-\mathrm{i}(\phi-\mathrm{i}\epsilon)})}{\Lambda(\mathrm{e}^{-\mathrm{i}(\phi-\mathrm{i}\epsilon)})}\right)\,\mathrm{d}r\mathrm{dA}\,\mathrm{d}\phi$
@@
$\times\left(2N\int_{SO(2N)}\Lambda(1)^{r}\mathrm{dA}-2\int_{SO(2N)}\mathrm{e}^{-\mathrm{i}(\phi-\mathrm{i}\epsilon)}\frac{\Lambda(1)^{r}\Lambda^{\prime}(\mathrm{e}^{-\mathrm{i}(\phi-\mathrm{i}\epsilon)})}{\Lambda(\mathrm{e}^{-\mathrm{i}(\phi-\mathrm{i}\epsilon)})}\mathrm{dA}\right)\,\,\mathrm{d}r\mathrm{d}\phi.$ (155)
@@
$\times\left(2N\left(1+\frac{r(r-1)^{2}}{2N}+\mathcal{O}(N^{-2})\right)+2\mathcal{U}(N,r,\mathrm{i}(\phi-\mathrm{i}\epsilon))\left(1+\mathcal{O}(N^{-2})\right)\right)\,\,\mathrm{d}r\mathrm{d}\phi.$ (156)
@@
$\times\left(2N\left(1+\frac{r(r-1)^{2}}{2N}\right)+2\mathcal{U}(N,r,\mathrm{i}\phi)\right)\left(1+\mathcal{O}\left(N^{-2}\right)\right)\,\,\mathrm{d}rf(\phi)\mathrm{d}\phi.$ (157)
@@
[…]
$\left.\left.\quad\quad\quad\quad\times\left(2N\left(1+\frac{r(r-1)^{2}}{2N}\right)+2\mathcal{U}(N,r,\mathrm{i}\phi)\right)\,\,\right|_{r=0}\right]f(\phi)\mathrm{d}\phi$
@@
$=\$$\frac{1}{2\pi}\int_{-\pi}^{\pi}\left[2N-2z(-2\mathrm{i}\phi)-2\mathrm{e}^{-2N\mathrm{i}\phi}z(2\mathrm{i}\phi)\right]f(\phi)\mathrm{d}\phi$
$=\$$\frac{1}{2\pi}\int_{-\pi}^{\pi}\left[2N-\frac{2}{1-\mathrm{e}^{2\mathrm{i}\phi}}-2\frac{\mathrm{e}^{-2N\mathrm{i}\phi}}{1-\mathrm{e}^{-2\mathrm{i}\phi}}\right]f(\phi)\mathrm{d}\phi$
$=\$$\frac{1}{2\pi}\int_{-\pi}^{\pi}\left[2N-\frac{2(\cos(\phi)-\mathrm{i}\sin(\phi))}{-2\mathrm{i}\sin(\phi)}+\right$.
@@
$=\$$\frac{1}{2\pi}\int_{-\pi}^{\pi}\left(2N-1+\sin\left((2N-1)\phi\right)/\sin(\phi)\right)f(\phi)\mathrm{d}\phi.$ (159)
@@
The $G$-function $G(2r+1)$ has zeros at $r\in\{-n/2\|n\in\mathbb{N}\}$, each one of multiplicity $n$. If we expand $G(2r+1)^{-1}$ as a Laurent series about n/2-n/2, then, we get terms proportional to $(r+n/2)^{-n},(r+n/2)^{-n+1},(r+n/2)^{-n+2},\dots$ etc. So, when we expand $G(2r+1)^{-1/2}$ around n/2-n/2, we get a Puiseaux series with terms proportional to $(r+n/2)^{-n/2},(r+n/2)^{-n/2+1},(r+n/2)^{-n/2+2}\dots$ etc. Note that we do not get powers of $(r+n/2)^{-n/2+1/2},\ (r+n/2)^{-n/2+3/2}$ etc., since this series must square to the Laurent series for $G(2r+1)^{-1}$.
We also know that $\Gamma(2r+1)$ has poles at $\{-n/2\|n\in\mathbb{N}\}$. In the Laurent expansion of $\Gamma(2r+1)$ around such a point n/2-n/2, there are terms proportional to $(r+n/2)^{-1},(r+n/2)^{0},(r+n/2)^{1}$ etc. This means that in the Puiseaux expansion of $\Gamma(2r+1)^{1/2}$ around the pole at n/2-n/2, we have terms proportional to $(r+n/2)^{-1/2},(r+n/2)^{1/2},(r+n/2)^{3/2}$ etc. Again, we do not get integer powers in this series, since it must square to the Laurent series for $\Gamma(2r+1)$.
Consider $\Gamma(2r+1)^{1/2}G(2r+1)^{-1/2}$ at a negative integer $\{-m\|m\in\mathbb{N}\}$. The series expansion of this product at m-m will have terms proportional to $(r+m)^{-m-1/2},(r+m)^{-m+1/2},\dots$ etc, from multiplying the Puiseaux series together. This series will not have a $(r+m)^{-1}$ term, so will not contribute a residue to (158).
@@
Figure 1: One level density for the excised $SO(24)$ ensemble with $\chi=\log(0.0001)$, generated numerically using $10^{7}$ matrices (histogram), with the first four terms (up to $k=3$) from (158) plotted as the smooth curve. Both plots are normalised to have unit area under the curve.